OpenStudy (dtan5457):

When expressed as a single term, what number is the numerical coefficient of the sqrt of 3? (Problem in comments)

OpenStudy (dtan5457):

\[\sqrt{75}-\frac{ 4 }{ 2\sqrt{3} }\]

OpenStudy (alexandervonhumboldt2):

@DanJS

OpenStudy (dtan5457):

@Nnesha

OpenStudy (danjs):

hi

OpenStudy (dtan5457):

yo.

OpenStudy (dtan5457):

i should rationalize this right?

OpenStudy (danjs):

what is the prob?

OpenStudy (dtan5457):

multiply square root of 75 by 2 sq rt of 3?

OpenStudy (dtan5457):

its up there ^^

OpenStudy (danjs):

\[\frac{ 2\sqrt{3} }{ 2\sqrt{3} } * [\frac{ \sqrt{75} }{ 1 } - \frac{ 4 }{ 2\sqrt{3} }]\]

OpenStudy (danjs):

multiply the whole thing by 1, or 2 root 3 over 2 root 3

OpenStudy (dtan5457):

just for radical 75 or 4/ 2 radical 3 as well?

OpenStudy (danjs):

distribute it to both of the terms

OpenStudy (dtan5457):

if i do just to one, the value becomes uneven?

OpenStudy (danjs):

it wont change the value, you are basically multiplying by 1

OpenStudy (danjs):

no, it just will be messier, i think you want to simplify it

OpenStudy (dtan5457):

then can i do just to radical 75?

OpenStudy (danjs):

you can do it to any of the terms you want, it is just 1, multiplying by 1, but if you do it to the root 75 term too, it will simplify to get rid of the root...

OpenStudy (dtan5457):

yeah that's what i did. i multipled radical 75/1 but 2 radical 3

OpenStudy (dtan5457):

which equates to 30?

OpenStudy (danjs):

You want to be able to put them over a common denominator, then you can rationalize the denominator

OpenStudy (dtan5457):

sq rt 225=15 times 2=30

OpenStudy (danjs):

yep, 30 over 2root 3

OpenStudy (dtan5457):

then i subtracted the 4

OpenStudy (dtan5457):

26/2 sq rt 3

OpenStudy (danjs):

yeah, now you can rationalize the denominator

OpenStudy (dtan5457):

i have to rationalize again?

OpenStudy (danjs):

\[\frac{ 26 }{ 2\sqrt{3} } = \frac{ 13 }{ \sqrt{3} } \]

OpenStudy (danjs):

yeah, you just put them over a common denominator in the first step, so you could subtract the two, now you can rationalize the denominator

OpenStudy (danjs):

13root3 over 3

OpenStudy (danjs):

\[\frac{ 13 }{ \sqrt{3} }*\frac{ \sqrt{3} }{ \sqrt{3} } = \frac{ 13\sqrt{3} }{ 3 }\]

OpenStudy (dtan5457):

what i did wrong was i thought \[\frac{ 26 }{ 2\sqrt{3} }=13\sqrt{3}\]

OpenStudy (danjs):

it said to express it as a single term, so you had to put them both over the common denominator, then rationalize the root in the denominator

OpenStudy (danjs):

yeah forgot that root 3 times root 3, leaves a 3 in the bottom

OpenStudy (danjs):

you have more ?

OpenStudy (dtan5457):

not right now, but can you tell me the rule in which my 26/2 sq rt 3 does NOT equal 13 sq root 3

OpenStudy (danjs):

ok

OpenStudy (danjs):

remember that the square root is the same as 'to the 1/2 power right?

OpenStudy (dtan5457):

yes.

OpenStudy (danjs):

26/2 = 13, so you see it is 13 over root 3 right?

OpenStudy (dtan5457):

i first thought it was 13 times root 3

OpenStudy (danjs):

\[\frac{ 26 }{ 2 \sqrt{3}} = \frac{ 13 }{ \sqrt{3} }\]

OpenStudy (danjs):

just divided by 2, both top and bottom

OpenStudy (dtan5457):

ooh. golden words right there, thanks.

OpenStudy (danjs):

now multiply like this , it is a ratio of 1 \[\frac{ 13 }{ \sqrt{3} }*\frac{ \sqrt{3} }{ \sqrt{3} } = \frac{ 13\sqrt{3} }{ \sqrt{3}\sqrt{3} }\]

OpenStudy (danjs):

you get that part?

OpenStudy (dtan5457):

Yeah, i get that part, the co efficient is 13/3

OpenStudy (danjs):

\[\sqrt{3}\sqrt{3} = 3^{1/2}*3^{1/2} = 3^{1/2 + 1/2} = 3^{1} = 3\]

OpenStudy (danjs):

when you square a square root, you are left with the quantity inside the root

OpenStudy (danjs):

so the final answer is \[\frac{ 13\sqrt{3} }{ 3 }~~~~or ~~~~\frac{ 13 }{ 3 }*\sqrt{3}\]

OpenStudy (danjs):

coefficient on root 3, 13/3

OpenStudy (danjs):

u understand all of it now? 1) put over common denominator 2) rationalize the denominator 3) simplify

OpenStudy (dtan5457):

yup. thanks.